# Training Hybrid Deep Quantum Neural Network for Efficient Reinforcement Learning

### v4

@@ 1 Introduction
In this work, we introduce hDQNN-TD3, a hybrid quantum neural network architecture that solves this bottleneck using a novel approach: a classical surrogate that we named qtDNN will approximate the PQC’s gradient updates on the backward pass. We demonstrate the hDQNN-TD3 model on the Humanoid-v4 reinforcement learning benchmark with better performance over the widely adopted classical SOTA counterparts. This paper is organized as follows: we will introduce the method that enables scalable training of hDQNN with classical DNNs around PQCs in Sec. 2. Following this, Sec. 3 will introduce hDQNN-TD3 for continuous variable RL tasks and show the experimental results from Humanoid-v4 demonstrating that hDQNN-TD3 outperforms comparable classical methods as well as publicly reported models using SOTA RL architectures from Zhao (2024, 2023a, 2023b). The last Sec. 4 will discuss outlooks to train significantly larger models.
@@ 3.1 Problem set-up and motivation
Moreover, Fig. 5 in Appendix A.4 details the data flow inside one update epoch, making explicit which parameters are optimized at which stage.

### v5

# Training Hybrid Deep Quantum Neural Networks for Efficient Reinforcement Learning

@@ 1 Introduction
In this work, we introduce hDQNN-TD3, a hybrid quantum neural network architecture that approaches this bottleneck using a novel approach: a classical surrogate that we named qtDNN will approximate the PQC’s gradient updates on the backward pass. We demonstrate the hDQNN-TD3 model on the Humanoid-v4 reinforcement learning benchmark with better performance over the widely adopted classical SOTA counterparts. This paper is organized as follows: we will introduce the method that enables scalable training of hDQNN with classical DNNs around PQCs in Sec. 2. Following this, Sec. 3 will introduce hDQNN-TD3 for continuous variable RL tasks and show the experimental results from Humanoid-v4 demonstrating that hDQNN-TD3 outperforms comparable classical methods as well as publicly reported models using SOTA RL architectures from Zhao (2024, 2023a, 2023b). The last Sec. 4 will discuss outlooks to train significantly larger models.
@@ 3.1 Problem set-up and motivation
Moreover, Fig. 6 in Appendix D.2 details the data flow inside one update epoch, making explicit which parameters are optimized at which stage.